OpenStudy (peterg2001):

help please will medal @AlexandervonHumboldt2 @beastieman21 @k_lynn @Directrix @DanJS @Jesstho.-. @triciaal @GhostGamerYT @dtan5457 @loganisawesome151 @ (1/4x)-^2=________/x^2

OpenStudy (peterg2001):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

a^-b=1/a^b

OpenStudy (peterg2001):

@esshotwired

OpenStudy (peterg2001):

sure?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (alexandervonhumboldt2):

1/(1/4x)^2

OpenStudy (alexandervonhumboldt2):

simplify now

OpenStudy (peterg2001):

so the answer??? didnt get it

OpenStudy (peterg2001):

@hhelpplzzzz

OpenStudy (peterg2001):

NEED DONE now HELP please WILL MEDAL OR THREE

OpenStudy (alexandervonhumboldt2):

1/1/16x^2=16/x^2

OpenStudy (alexandervonhumboldt2):

medal at least an testimonilal because i shouldn't give a direct answer

OpenStudy (peterg2001):

sorry wrong questoin

OpenStudy (alexandervonhumboldt2):

ok

OpenStudy (dtan5457):